A 1250 kg car is pulling a 325 kg trailer. Together, the car and trailer have an acceleration of 2.15 m/sec^2. Determine the net force on the car-trailer combination. What is the net force on the trailer?
1
Expert's answer
2010-11-19T10:41:53-0500
The net force would be:
F = (M+m)a = (1250+325)*2.15 = 3386.25 N
F(trailer) =ma = 325*2.15 = 698.75 N